OpenStudy (here_to_help15):

I WILL FAN ANYONE WHO COMES IN!!!SOLVE 1+1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000

OpenStudy (anonymous):

1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000001

OpenStudy (here_to_help15):

Sorry satellite i will close this :(

OpenStudy (anonymous):

17 :)

OpenStudy (anonymous):

42

OpenStudy (anonymous):

1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000001

OpenStudy (anonymous):

lawl

OpenStudy (here_to_help15):

@damien

OpenStudy (here_to_help15):

@Deadpool25 , @Donblue22 , @sunnycali , @Firejay5 , @Joel_the_boss

OpenStudy (anonymous):

Did i win ???

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

@Skyz @stefrheart @Nnesha @kmullis6 @OswinOswald @TheSmartOne @ilovebmth1234

OpenStudy (firejay5):

1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000001

JoelTheBoss (joel_the_boss):

You guys are doing it all wrong! 1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000+1= 15-11%/10= 4

OpenStudy (anonymous):

i won and got it correct!

OpenStudy (here_to_help15):

lol

OpenStudy (anonymous):

look carefully and thats it

OpenStudy (here_to_help15):

@baby_boo @charlotte123 @CrazyCountryGirl @gibbs

OpenStudy (here_to_help15):

@JakeCollege

OpenStudy (here_to_help15):

@DolphinDawg

OpenStudy (stefrheart):

sig figs- significant figures

OpenStudy (stefrheart):

sorry i mean 1

OpenStudy (here_to_help15):

Can 2 ppl fan me?

OpenStudy (stefrheart):

1 sig fig

OpenStudy (here_to_help15):

yup @stefrheart

OpenStudy (gibbs):

what am i doing

OpenStudy (stefrheart):

lol my teacher did that to us on the first day of school but it was 20+1

OpenStudy (here_to_help15):

lol

OpenStudy (anonymous):

1000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000000001

OpenStudy (stefrheart):

It was a cruel joke

OpenStudy (here_to_help15):

@love10129151 @lil_yadira14 @Librarian

OpenStudy (here_to_help15):

@pondamo @Perce

OpenStudy (anonymous):

I swallow noodles :)

OpenStudy (gibbs):

can't help

OpenStudy (anonymous):

You should be making the walking dead SON

OpenStudy (stefrheart):

Hahaha!

OpenStudy (anonymous):

can't wait tho

OpenStudy (anonymous):

Hey wheres my fan?! XD

OpenStudy (anonymous):

someone answered before u

OpenStudy (anonymous):

but i will fan

OpenStudy (anonymous):

thanks :3

OpenStudy (anonymous):



OpenStudy (anonymous):

That one number that is the answer to your problem there, solved it. >.>

OpenStudy (anonymous):

Guys it's obviously 21